Aratox:

Identify the Inverse Function of g(x)=(1/2)x+2.

somnium:

hdbhdbh i forgor how to do those

somnium:

gimme a min

Aratox:

It's IB Precalculus, don't answer it if you haven't taken Precalc yet 😂

somnium:

shush i can do math even if i havent learned it yet, ik my capabilities, i just dont remember them

0bviouslyNotRakita:

g-1(x)= 1/2x-1 I think this is the correct answer

somnium:

if it's just interchanging variables and simplifying like i remember, then it should be like this \[g ^{-1}(x)=2x-4\]

somnium:

WOW THATS MY FIRST TIME USING THE EQUATION THINGY AND I DID IT RIGHT :noice:

Aratox:

@0bviouslynotrakita wrote:
g-1(x)= 1/2x-1 I think this is the correct answer
I believe that the answer was g^(-1)(x)=(1/2)x, how'd you get the -1 at the end of your equation?

Aratox:

At least, I got my g^(-1)(x)=(1/2)x, I don't have the correct answer...

0bviouslyNotRakita:

@aratox wrote:
@0bviouslynotrakita wrote:
g-1(x)= 1/2x-1 I think this is the correct answer
I believe that the answer was g^(-1)(x)=(1/2)x, how'd you get the -1 at the end of your equation?
By dividing the -2 to the positive 2. Thats how I got -1

Aratox:

@0bviouslynotrakita wrote:
@aratox wrote:
@0bviouslynotrakita wrote:
g-1(x)= 1/2x-1 I think this is the correct answer
I believe that the answer was g^(-1)(x)=(1/2)x, how'd you get the -1 at the end of your equation?
By dividing the -2 to the positive 2. Thats how I got -1
OHH that's what I missed... thank you :D

0bviouslyNotRakita:

@aratox wrote:
@0bviouslynotrakita wrote:
@aratox wrote:
@0bviouslynotrakita wrote:
g-1(x)= 1/2x-1 I think this is the correct answer
I believe that the answer was g^(-1)(x)=(1/2)x, how'd you get the -1 at the end of your equation?
By dividing the -2 to the positive 2. Thats how I got -1
OHH that's what I missed... thank you :D
your welcome :)